# CLASP: Adaptive Spectral Clustering for Unsupervised Per-Image Segmentation

### v1

@@ Abstract
We introduce CLASP (Clustering via Adaptive Spectral Processing), a lightweight framework for unsupervised image segmentation that operates without any labeled data or fine-tuning. CLASP first extracts per-patch features using a self-supervised ViT encoder (DINO); then, it builds an affinity matrix and applies spectral clustering. To avoid manual tuning, we select the segment count automatically with a eigengap-silhouette search, and we sharpen the boundaries with a fully connected DenseCRF. Despite its simplicity and training-free nature, CLASP attains competitive mIoU and pixel-accuracy on COCO-Stuff and ADE20K, matching recent unsupervised baselines. The zero-training design makes CLASP a strong, easily reproducible baseline for large unannotated corpora—especially common in digital advertising and marketing workflows such as brand-safety screening, creative asset curation, and social-media content moderation.
@@ 8 Acknowledgements
We thank our colleagues for their insightful feedback. This work was supported by Integral Ad Science. Special thanks to Shirley Davila for allowing us to photograph her pets—Petey, Stevie, Minnie, and Annie—who served as models in Fig. 4.

### v2

@@ Abstract
We introduce CLASP (Clustering via Adaptive Spectral Processing), a lightweight framework for unsupervised image segmentation that operates without any labeled data or finetuning. CLASP first extracts per patch features using a self supervised ViT encoder (DINO); then, it builds an affinity matrix and applies spectral clustering. To avoid manual tuning, we select the segment count automatically with a eigengap silhouette search, and we sharpen the boundaries with a fully connected DenseCRF. Despite its simplicity and training free nature, CLASP attains competitive mIoU and pixel accuracy on COCO Stuff and ADE20K, matching recent unsupervised baselines. The zero training design makes CLASP a strong, easily reproducible baseline for large unannotated corpora especially common in digital advertising and marketing workflows such as brand safety screening, creative asset curation, and social media content moderation.
@@ 8 Acknowledgements
We thank our colleagues for their insightful feedback. This work was supported by Integral Ad Science. We also thank Victor Rambaud for the many meaningful conversations that helped shape this work. Special thanks to Shirley Davila for allowing us to photograph her pets—Petey, Stevie, Minnie, and Annie—who served as models in Fig. 4.